OpenStudy (anonymous):

The steps to derive the quadratic formula are shown below: Step 1: ax2 + bx + c = 0 Step 2: x2 + b over ax + c over a = 0 Step 3: x2 + b over a x = - c over a Step 4: x2 +b over a x + square of quantity of b over 2 times a = -c over a + square of quantity of b over 2 times a Step 5: x plus b over 2 times a, all squared = -c over a + square of quantity of b over 2 times a Step 6: x + b over 2 times a = square root of all of negative c over a plus the square of the quantity b over 2 times a Step 7: x = - b over 2 times a ± square root of all of negative c over a plus the square of

OpenStudy (anonymous):

Step 7: x = - b over 2 times a ± square root of all of negative c over a plus the square of the quantity b over 4 times a = -b over 2 times a ± square root of the quantity of b squared minus 4 times a times c over 4 times a squared= negative b plus or minus square root of b squared minus 4 times a times c over 2 times a Which of the following statements best describes the first incorrect step? In Step 4, square of quantity of b over 2 times a instead of b over 2 times a was added on both sides. In Step 5, x plus b over 2 times a, all squared was written instead of x plus b over 4 times a, all squared . In Step 6, -c over a was written instead of c over a inside the root. In Step 7, b over 4 times a, all squared was written instead of square of quantity of b over 2 times a inside the root.

OpenStudy (anonymous):

I believe step 4 and 5 are right but i dont know about 6 and 7

OpenStudy (anonymous):

@Hero @SolomonZelman @Abhisar @Ail.A @JReed231 @Nurali @uri @kropot72 @kirbykirby @KlOwNlOvE @xapproachesinfinity @NoCopsForMiles

OpenStudy (anonymous):

someone please i will give medal

OpenStudy (anonymous):

@tejasvir @ganeshie8 @Gin_Ichimaru @geekfromthefutur @GoldenEagle @graffiboy1 @BassCatcher15 @Blink-182 @bibby @BVBLegacy @Brainybeauty @dan815 @dynasty<> @tejasvir @TheMilkyWay @miraclemilktea2 @math92130 @iceicebaby

OpenStudy (anonymous):

sorry I cant give you the answer otherwise I will get banned and we don't want this to happen

OpenStudy (anonymous):

can you help?

OpenStudy (anonymous):

@graffiboy1

OpenStudy (anonymous):

the quadratic formula is

OpenStudy (anonymous):

\[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[ax ^{2}+bx+c=0\]

OpenStudy (anonymous):

\[x=\frac{ -1 \pm \sqrt{1^{2}-4(1)(1)} }{ 2(1) }\]

OpenStudy (anonymous):

\[x=\frac{ -1 \pm \sqrt{1-4} }{ 2 }\]

OpenStudy (anonymous):

\[x=\frac{ -1 \pm \sqrt{-3} }{ 2 }\]

OpenStudy (anonymous):

\[x=\frac{ \pm \sqrt{-4} }{ 2 }\]

OpenStudy (anonymous):

\[x=\frac{ \pm 2 }{ 2 }\]

OpenStudy (anonymous):

\[x=\pm1\]

OpenStudy (anonymous):

are you sure bc you cant add -1 and \[\sqrt{-3}\]

OpenStudy (anonymous):

@phi can you explain this?

OpenStudy (phi):

owl-lover, your post is hard to read when typed out in words.

OpenStudy (anonymous):

@owl-lover

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

is step 7 the answer? here:\[\frac{ -b }{ 2a }\pm \sqrt{\frac{ -c }{ a }+(\frac{ b }{ 4a })^2}=\frac{ -b }{ 2a }\]

OpenStudy (phi):

but the last step might be the one with an error

OpenStudy (anonymous):

nvm this is really hard thats not all of 7

OpenStudy (anonymous):

can you screenshot the question

OpenStudy (phi):

the (b/4a)^2 in step 7 should read (b/2a)^2

OpenStudy (anonymous):

not sure how

OpenStudy (anonymous):

and yes thats what i was thinking

OpenStudy (anonymous):

its not working but i think i got it could you help with another